ROTATIONAL SPECTRUM SPECTRUM AND COUPLED-CLUSTER CALCULATIONS OF SILICON OXY-SULFIDE, O=Si=S

\underline{S. THORWIRTH}, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; L. A. MÜCK, J. GAUSS, Institut für Physikalische Chemie, Universität Mainz, 55099 Mainz, Germany; F. TAMASSIA, Dipartimento di Chimica Fisica e Inorganica, Universitá di Bologna, I-40136 Bologna, Italy; V. LATTANZI, M. C. McCARTHY, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, and School of Engineering and Applied Science, Harvard University, Cambridge, MA 02138.

Silicon oxysulfide, OSiS, and seven of its minor isotopic species have been characterized for the first time in the gas phase at high spectral resolution by means of Fourier-transform microwave spectroscopy. The equilibrium structure of OSiS has been determined from a combination of experimental ground state rotational constants and calculated vibrational corrections to those. The structural parameters are in good agreement with values from high-level quantum-chemical calculations using coupled-cluster techniques together with sophisticated additivity and extrapolation schemes.